OpenStudy (anonymous):

Patrick and Brooklyn are making decisions about their bank accounts. Patrick wants to deposit $300 as a principle amount, with an interest of 3% compounded quarterly. Brooklyn wants to deposit $300 as the principle amount, with an interest of 5% compounded monthly. Explain which method results in more money after 2 years. Show all work.

OpenStudy (popeyes):

Patrick: P = $300 I = 3% compounded quarterly T = 2 years Brooklyn: P = $300 I = 5% compounded monthly T = 2 years Patrick: amount = \[p(1 + \frac{ r }{ 100 })^n\] \[CI = 300 (1 + \frac{ 3 }{ 400 }) = $318.48\] Brooklyn: amount = \[300 (1 + \frac{ 5 }{ 1200 })^{24} = $331.48\] Compounded monthly is more beneficial than compounded quarterly.